You place $25,000 in a savings account paying annual compound interest of 8% for three years and then move it into a savings account that pays 10% interest compounded annually. How much will your money have grown at the end of six years?




1
Expert's answer
2021-03-08T07:12:28-0500

Amount at end of first three years

"A=P\times(1+r)^t"

Where, A= Total amount at year end, P = Principal, r= rate per annum and t = time in years

"A=25,000\times(1.08)^3=31,492.80"

Amount at end of next 3 years

"A= 31,492.8\times(1.1)^3=41,916.92"

At end of 6 years, the money would have grown to $41,916.92